OpenStudy (highschoolmom2010):

Find the area of the regular polygon. Give the answer to the nearest tenth. decagon with a side of 4 cm 123.1 cm^2 129.4 cm ^2 246.2 cm ^2 139.8 cm ^2

OpenStudy (jdoe0001):

 how many degrees do you think each of the 6 sides will make in that hexagon?

OpenStudy (highschoolmom2010):

no clue what does a hexagon have to do with my question jw

OpenStudy (jdoe0001):

hmmm ohhh..ahemm . it's a decagon lol, I thought it was a hexagon hehe

OpenStudy (highschoolmom2010):

lol understandable

OpenStudy (highschoolmom2010):

thanks for trying

OpenStudy (jdoe0001):

well... ahemm heheh

OpenStudy (jdoe0001):

trying to make a quick graph of the central angle =)

OpenStudy (highschoolmom2010):

ok lol

OpenStudy (highschoolmom2010):

anyone.....

OpenStudy (jdoe0001):

heheh,darn it's taking longer than expected =)

OpenStudy (rational):

Area of regular polygon = \(\large \frac{n}{4}\times s^2\cot(\frac{180}{n})\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (rational):

for decagon \(n = 10\) \(s = 4\) plug into your calculator. do you need a derivation for the formula ?

OpenStudy (rational):